OpenStudy (anonymous):

MEDAL::What is the simplified form of the seventh root of x to the fifth power times the seventh root of x to the fifth power? how do i solve this???

OpenStudy (anonymous):

\[\sqrt[7]{x^5}\sqrt[7]{x^5}\]?

OpenStudy (anonymous):

first step is to write \[\sqrt[7]{x^5\times x^5}\] adding exponents gives \[\sqrt[7]{x^{10}}\] but this is not in simplest radical form since \(10>7\)

OpenStudy (anonymous):

7 goes in to ten once, with a remainder of three so \[\sqrt[7]{x^{10}}=x\sqrt[7]{x^3}\]